# Evolving Beyond Snapshots: Harmonizing Structure and Sequence via Entity State Tuning for Temporal Knowledge Graph Forecasting

### 12 Mar 2026 v2

@@ Abstract
Temporal knowledge graph (TKG) forecasting requires predicting future facts by jointly modeling structural dependencies within each snapshot and temporal evolution across snapshots. However, most existing methods are stateless: they recompute entity representations at each timestamp from a limited query window, leading to episodic amnesia and rapid decay of long-term dependencies. To address this limitation, we propose Entity State Tuning (EST), an encoder-agnostic framework that endows TKG forecasters with persistent and continuously evolving entity states. EST maintains a global state buffer and progressively aligns structural evidence with sequential signals via a closed-loop design. Specifically, a topology-aware state perceiver first injects entity-state priors into structural encoding. Then, a unified temporal context module aggregates the state-enhanced events with a pluggable sequence backbone. Subsequently, a dual-track evolution mechanism writes the updated context back to the global entity state memory, balancing plasticity against stability. Experiments on multiple benchmarks show that EST consistently improves diverse backbones and achieves state-of-the-art performance, highlighting the importance of state persistence for long-horizon TKG forecasting. The code is published at https://github.com/yuanwuyuan9/Evolving-Beyond-Snapshots.
@@
[…]
@@ 1 Introduction
Temporal Knowledge Graphs (TKGs) extend static facts into quadruples $(s,r,o,t)$, organizing knowledge as a sequence of chronological snapshots Cai et al. (2024a); Wang et al. (2024a); Liu et al. (2025a). This temporal grounding is pivotal for reasoning in dynamic environments, enabling applications ranging from Question Answering Wang et al. (2024b) to Recommender Systems Wang et al. (2025).
@@
Existing approaches typically follow a divide-and-conquer strategy, biasing toward either structural modeling, represented by RE-GCNLi et al. (2021), or sequential linearization, as exemplified by TiSASRecLi et al. (2020). Despite architectural diversity, existing TKG methods largely operate under a stateless reasoning paradigm. As illustrated in Figure 1(a), entity representations are independently reconstructed from each snapshot, preventing historical information from being persistently accumulated across time. As a result, long-range dependencies are gradually diluted as temporal reasoning unfolds, a phenomenon we term episodic amnesia. In contrast, Figure 1(b) highlights the necessity of modeling entities as persistent state carriers to naturally bridge structural and temporal dynamics.
@@ 4.2 Unified Temporal Context Encoder
We formulate EST as a meta-framework capable of empowering diverse sequential architectures. Instead of coupling with a specific encoder, we abstract sequence modeling as a mapping from the stream $\mathbf{U}=[\mathbf{u}_{1},\dots,\mathbf{u}_{L}]$ to contextualized hidden states. For the $k$-th step, the update is defined as:
$\mathbf{h}_{k}=\mathcal{T}\big(\mathbf{h}_{<k},\mathbf{u}_{k};\theta_{\text{seq}}\big),\quad\mathbf{y}_{k}=\mathcal{P}(\mathbf{h}_{k}),$ (6)
where $\mathcal{T}(\cdot)$ represents a generalized state transition function (e.g., RNN or Mamba blocks) and $\mathcal{P}(\cdot)$ is a projection head. We validate this extensibility by benchmarking various backbones in Section 5.2.
@@
$\alpha_{k}\propto\exp\!\left(\mathbf{v}^{\top}\tanh\!\left(\mathbf{W}_{\text{att}}\left[\mathbf{y}_{k};\,\mathbf{e}_{r};\,\mathbf{e}_{t}\right]\right)\right).$ (7)
@@ 4.3 De-confounded State Evolution
Slow System (Consolidated Memory)
@@
Table 1: Overall performance on four benchmark datasets. Best results are in bold, second best are underlined
@@ 5.2 Overall Performance Comparison: RQ1
Efficiency and variant selection.
EST adopts a plug-and-play sequential architecture, abstracting the temporal evolution into a generalized state transition function $\mathcal{T}$. We instantiate this module with four backbones. The specific mathematical formulations and update dynamics for each variant are detailed in Appendix D. We then benchmark them on ICEWS14 in terms of convergence, model size, and wall-clock training time.
As illustrated in Figure 2(a), all variants converge rapidly and stably, surpassing 0.50 test MRR within a few epochs, suggesting that the gains primarily come from EST’s stateful reasoning rather than a specific sequential backbone. Among them, EST-Transformer achieves the best final performance, whereas EST-Mamba provides a strong and smooth alternative. Efficiency-wise, Table 2 shows that EST-Transformer trains fastest (18.58 minutes at epoch 20), with EST-Mamba close behind (19.81 minutes), both outperforming recurrent counterparts. Moreover, EST-Mamba is the most parameter-efficient in Figure 2(b) (2.08 MB vs. 2.53 MB for EST-Transformer) while maintaining competitive metrics in Table 1.
Accordingly, we use EST-Transformer as the default (accuracy-oriented) instantiation and EST-Mamba as a lightweight counterpart for subsequent analyses; unless stated otherwise, later experiments report results for these two variants.
5.3 Ablation Study: RQ2
@@ 5.3.1 Key Modules Ablation Study
(a) ICEWS14 Performance Trends.
(b) ICEWS18 Performance Trends.
Figure 7: Assessing temporal generalization in EST-Transformer via chronological data truncation.
@@ 5.5 Temporal Generalization: RQ4
To evaluate robustness against distribution shifts, we employ a Chronological Data Truncation protocol, training models solely on the earliest $X\%$ of history ($X\in\{50,\dots,90\}$).
Figure 7 illustrates the resilience of EST-Transformer (consistent trends for EST-Mamba are provided in Appendix E). Despite expected attenuation under sparse supervision, EST maintains a robust lead over its stateless ablation across all levels. Most strikingly, EST demonstrates exceptional data efficiency: even when restricted to only 50% of the training history, EST-Transformer achieves an MRR of 0.416 on ICEWS18, significantly outperforming fully-supervised SOTA baselines like DiffuTKG (0.367). This confirms that our state-centric paradigm captures intrinsic evolutionary dynamics more effectively than stateless approaches.
@@ 6 Conclusion
First, EST maintains global entity state buffers, incurring memory overhead that scales linearly with the number of entities ($\mathcal{O}(|\mathcal{E}|d)$). While efficient for common benchmarks, this design may become costly for web-scale graphs or deployments with strict memory budgets.
Second, our closed-loop evolution assumes a temporally ordered processing regime and a relatively stable entity universe. Handling previously unseen entities or highly dynamic entity sets would require additional mechanisms (e.g., state initialization or consolidation strategies), which we leave for future work.
Finally, our evaluation focuses on four widely used benchmarks under standard chronological splits. Although these settings cover diverse dynamics, real-world TKGs may exhibit stronger non-stationarity, exogenous shocks, or richer supervision signals (e.g., attributes and text), which are not explicitly modeled in the current framework.
@@ Appendix A Datasets
(a) ICEWS14 Performance Trends.
(b) ICEWS18 Performance Trends.
Figure 8: Assessing temporal generalization in EST-Mamba via chronological data truncation.
@@ Appendix D Sequential Backbones
Appendix E Temporal Generalization
Complementing the analysis of EST-Transformer in Section 5.5, we present the corresponding temporal generalization trajectories for the EST-Mamba variant in Figure 8. This analysis aims to verify whether the data efficiency and robustness gains are intrinsic to the EST framework itself or dependent on specific attention mechanisms.
Universal Robustness and Efficiency.
[…]

### 16 Apr 2026 v3

@@ Abstract
Temporal knowledge graph (TKG) forecasting requires predicting future facts by jointly modeling structural dependencies within each snapshot and temporal evolution across snapshots. However, most existing methods are stateless: they recompute entity representations at each timestamp from a limited query window, leading to episodic amnesia and rapid decay of long-term dependencies. To address this limitation, we propose Entity State Tuning (EST), an encoder-agnostic framework that endows TKG forecasters with persistent and continuously evolving entity states. EST maintains a global state buffer and progressively aligns structural evidence with sequential signals via a closed-loop design. Specifically, a topology-aware state perceiver first injects entity-state priors into structural encoding. Then, a unified temporal context module aggregates the state-enhanced events with a pluggable sequence backbone. Subsequently, a dual-track evolution mechanism writes the updated context back to the global entity state memory, balancing plasticity against stability. Experiments on multiple benchmarks show that EST consistently improves diverse backbones and achieves state-of-the-art performance, highlighting the importance of state persistence for long-horizon TKG forecasting.
@@
Siyuan Li1thanks: Corresponding authors.thanks: Equal contribution., Yunjia Wu122footnotemark: 2, Yiyong Xiao [^2] footnotemark: 2, Pingyang Huang1, Peize Li3, Ruitong Liu411footnotemark: 1, Yan Wen5, Te Sun6 1 Dalian University of Technology  2 Shenzhen University of Advanced Technology 3 King’s College London  4 Peking University 5 Beijing Institute of Technology  6 Shanghai Jiao Tong University yuanlsy@mail.dlut.edu.cn, ruitong.jerry@gmail.com
@@ 1 Introduction
Temporal Knowledge Graphs (TKGs) extend static facts into quadruples $(s,r,o,t)$, organizing knowledge as a sequence of chronological snapshots Cai et al. (2024a); Wang et al. (2024a); Liu et al. (2025a); Lin et al. (2026). This temporal grounding is pivotal for reasoning in dynamic environments, enabling applications ranging from Question Answering Wang et al. (2024b); Liu et al. (2026) to Recommender Systems Wang et al. (2025).
@@
Existing TKG forecasting methods typically follow a divide-and-conquer paradigm, focusing either on structural dependency modeling, exemplified by RE-GCN Li et al. (2021), or on sequential linearization, as in TiSASRec Li et al. (2020). Despite their architectural diversity, most of these methods still operate under a fundamentally stateless paradigm. As shown in Figure 1(a), entity representations are reconstructed independently at each snapshot, preventing historical evidence from being persistently accumulated over time. As temporal reasoning proceeds, long-range dependencies are therefore progressively weakened, leading to what we term episodic amnesia. By contrast, Figure 1(b) highlights the necessity of modeling entities as persistent state carriers to naturally bridge structural and temporal dynamics.
@@ 4.2 Unified Temporal Context Encoder
We formulate EST as an encoder-agnostic framework that supports diverse sequential architectures within the same state-tuning pipeline. Given the stream $\mathbf{U}=[\mathbf{u}_{1},\dots,\mathbf{u}_{L}]$, the sequence backbone maps each causal prefix to a contextualized representation:
$\mathbf{h}_{k}=\mathcal{T}\big(\mathbf{u}_{\leq k};\theta_{\text{seq}}\big),\quad\mathbf{y}_{k}=\mathcal{P}(\mathbf{h}_{k}),$ (6)
where $\mathbf{u}_{\leq k}=[\mathbf{u}_{1},\dots,\mathbf{u}_{k}]$ denotes the causally visible prefix up to step $k$, $\mathbf{h}_{k}$ is the contextualized representation produced by the chosen backbone at position $k$, and $\mathcal{P}(\cdot)$ is a projection head. For notational uniformity, we use the same form for all backbone instantiations, while their concrete formulations (e.g., RNN, LSTM, Transformer, and Mamba) are provided in Appendix D. We validate this flexibility empirically in Section 5.2.
@@
$\alpha_{k}\propto\exp\!\left(\mathbf{v}^{\top}\tanh\!\left(\mathbf{W}_{\text{att}}\left[\mathbf{y}_{k};\,\mathbf{e}_{r};\,\varphi(\Delta t)\right]\right)\right).$ (7)
@@ 4.3 De-confounded State Evolution
Slow System (Consolidated Memory).
@@
Table 1: Overall performance on four benchmark datasets. Best results are in bold, second best are underlined.
@@ 5.1 Experiment Settings
Table 3: Additional memory introduced by the two FP32 global state buffers in EST (d=64d{=}64).
Dataset #Entities Buffer memory (FP32)
ICEWS14 7,128 3.48 MB
ICEWS18 10,094 4.93 MB
ICEWS05-15 23,033 11.25 MB
GDELT 7,691 3.76 MB
Extrapolation 1,000,000 488.28 MB ($\sim$0.49 GB)
Table 4: Measured runtime and peak-memory overhead of EST on ICEWS14, compared with a matched stateless ablation (w/o State).
Setting s/epoch ms/step Peak MB
EST-Mamba 60.77 105.09 173.80
w/o State 59.60 102.46 165.21
$\Delta$ +1.96% +2.57% +5.20%
EST-Transformer 61.51 105.22 111.61
w/o State 61.04 104.96 101.12
$\Delta$ +0.77% +0.25% +10.37%
@@ 5.2 Overall Performance Comparison: RQ1
5.2.1 Efficiency and Scalability
EST is encoder-agnostic, abstracting temporal evolution as a generalized state transition function $\mathcal{T}$. We instantiate $\mathcal{T}$ with four backbones; the corresponding formulations are provided in Appendix D. As shown in Figure 2 and Table 2, all variants converge reliably, indicating that the gains primarily come from persistent state modeling rather than a particular sequence encoder. Among them, EST-Transformer achieves the best overall accuracy, while EST-Mamba offers the most parameter-efficient alternative with competitive performance. Accordingly, we use EST-Transformer as the default accuracy-oriented instantiation and EST-Mamba as a lightweight counterpart in subsequent analyses.
Beyond encoder-level efficiency, we further analyze the overhead introduced by persistent global states. EST maintains two non-trainable entity buffers, $\mathbf{S}_{\mathrm{fast}}$ and $\mathbf{S}_{\mathrm{slow}}$, each of size $|\mathcal{E}|\times d$. The resulting memory overhead is therefore linear in the number of entities, i.e., $\mathcal{O}(|\mathcal{E}|d)$. Under our default setting (d=64d{=}64, FP32), the total buffer footprint is $2\times|\mathcal{E}|\times 64\times 4=512|\mathcal{E}|\text{ bytes},$ which remains at MB scale on standard benchmarks (Table 3).
Importantly, EST does not scan all entity states at each step. State access is sparse: for each batch, we only read and update entities appearing in the current history window. As a result, the runtime remains dominated by the sequential backbone and local structural encoding, while the global buffers introduce only lightweight indexing and in-place updates. Accordingly, the overall sequence modeling complexity remains that of the chosen backbone, namely $\mathcal{O}(Ld^{2})$ for RNN/LSTM, $\mathcal{O}(L^{2}d)$ for Transformer, and $\mathcal{O}(Ld)$ for Mamba.
To quantify the practical overhead of persistent states, we compare EST against a strictly matched stateless ablation (w/o State) on ICEWS14. Table 4 shows that the additional runtime is modest for both backbones: for Mamba, EST introduces only $1.96\%$ s/epoch and $2.57\%$ ms/step, while for Transformer the increase is $0.77\%$ and $0.25\%$, respectively. The peak-memory increase is also limited ($5.20\%$ for Mamba and $10.37\%$ for Transformer). Overall, EST introduces a linear memory overhead in $|\mathcal{E}|$ and only minor runtime cost in practice, making the global-buffer design practical on standard TKGs and viable at larger scales.
@@ 5.4 Entity State Evolution Analysis: RQ3
(a) EST-Transformer on ICEWS14.
(b) EST-Transformer on ICEWS18.
(c) EST-Mamba on ICEWS14.
(d) EST-Mamba on ICEWS18.
Figure 7: Temporal generalization under chronological data truncation. The first row shows EST-Transformer and the second row shows EST-Mamba; the left and right columns correspond to ICEWS14 and ICEWS18, respectively. EST consistently maintains a clear advantage over its stateless ablation across different training-history budgets.
@@ 5.5 Temporal Generalization: RQ4
To evaluate robustness under temporal distribution shift, we adopt a Chronological Data Truncation protocol, where models are trained only on the earliest $X\%$ of the training history ($X\in\{50,60,70,80,90\}$).
Figure 7 shows that EST consistently outperforms its stateless counterpart across all truncation levels on both ICEWS14 and ICEWS18, for both Transformer and Mamba backbones. Although performance degrades as less history is observed, the advantage of persistent states remains stable, indicating that the benefit of EST is not tied to a particular sequence encoder but is intrinsic to the state-centric paradigm.
Notably, EST also exhibits strong data efficiency. On ICEWS18 with only 50% of the training history, EST-Transformer still achieves an MRR of 0.416, surpassing fully-supervised baselines such as DiffuTKG (0.367), while EST-Mamba reaches 0.401, also exceeding DiffuTKG and CognTKE (0.352). These results suggest that persistent entity states capture transferable evolutionary dynamics beyond the observed window, leading to stronger temporal generalization than stateless modeling.
@@ 6 Conclusion
First, it maintains global entity state buffers, introducing a memory overhead that scales linearly with the number of entities, i.e., $\mathcal{O}(|\mathcal{E}|d)$; while practical on standard benchmarks, this may become costly for web-scale graphs or memory-constrained deployments. Second, EST assumes temporally ordered processing and a relatively stable entity universe, and thus does not explicitly address previously unseen entities or highly dynamic entity sets. Finally, our evaluation is limited to four standard benchmarks under chronological splits; real-world TKGs may exhibit stronger non-stationarity, exogenous shocks, or richer supervision signals (e.g., attributes and text), which are beyond the scope of the current study.
@@ Appendix D Sequential Backbones
Table 9: Worked example of history retrieval and query-specific attention on ICEWS14. For brevity, we show the most attended events in the retrieved history window.
Order Historical event $\Delta t$ Attention
1 (Refugee_(Afghanistan), Make_a_visit, Iran, 7944) 72 0.1753
2 (Refugee_(Afghanistan), Make_a_visit, Iran, 4728) 3288 0.1403
3 (Refugee_(Afghanistan), Make_a_visit, Iran, 5304) 2712 0.1412
4 (Refugee_(Afghanistan), Make_a_visit, Iran, 4848) 3168 0.1380
5 (Refugee_(Afghanistan), Make_a_visit, Iran, 4992) 3024 0.1344
@@
Appendix E End-to-End Workflow on ICEWS14
To make the formalization in Sections 34 more concrete, we present a worked example from ICEWS14 that instantiates Algorithm 1 on a real test query.
Step 1. Query formulation. Consider the test query:
$(\texttt{Refugee\_(Afghanistan)},\texttt{Make\_a\_visit},?,8016),$}
whose gold tail is Iran. EST treats this as a temporal forecasting query and retrieves the history of the head entity before timestamp $t=8016$.
Step 2. History retrieval and attention. Table 9 lists the most attended events in the retrieved history window. The history shows a repeated interaction pattern with Iran under the relation Make_a_visit, and the most recent event receives the largest attention weight, consistent with temporal forecasting intuition.
Step 3. Forward scoring. Each retrieved event is encoded into a state-aware structural representation and then passed through the sequence backbone to produce contextualized hidden states. Query-specific attention aggregates them into a context vector $\mathbf{c}_{s}$, which is used to score candidate tails via $\phi(\mathbf{c}_{s},r,o)$. In this example, repeated visits to Iran provide the dominant evidence for ranking the gold tail.
Step 4. Persistent-state write-back. After the forward pass, EST writes the resulting context back to the persistent states of the head entity through the fast and slow buffers. This closed-loop update allows future predictions involving Refugee_(Afghanistan) to depend on accumulated temporal evidence rather than being reconstructed independently at each snapshot.
Takeaway. This example illustrates how EST processes a real query end-to-end: query formulation, history retrieval, query-aware aggregation, tail scoring, and persistent state evolution.